# Temperature-Noise Interplay in a Coupled Model of Opinion Dynamics

### v2

@@ Abstract
We consider a coupled system mimicking opinion formation under the influence of a group of $q$ neighbors ($q$-lobby) that consists of an Ising part governed by temperature-like parameter $T$ and a voter dynamics parameterized by noise probability $p$ (independence of choice). Using rigorous analytical calculations backed by extensive Monte Carlo simulations, we examine the interplay between these two quantities. Based on the theory of phase transitions, we derive the relation between $T$ and $p$ at the critical line dividing the ordered and disordered phases, which takes a very simple and generic form $T(p-a)=b$ in the high temperature limit. For specific lobby sizes, we show where the temperature are noise are balanced, and we hint that for large $q$, the temperature-like dynamics prevails.
@@ I Introduction
By coupling these two generic dynamics we examine the conditions that are necessary for one of the models to prevail, thus finding the relation between the temperature and noise (independence). Additionally, owing to the parametrization of the model with the lobby size $q$, we show how the number of active neighbors affects the above relation and alters the system’s behavior.
@@ II Model description and its fundamental properties
The system consists of $N$ agents, each holding a binary opinion $s_{i}=\pm 1$, acting on top of a complete graph, described by total magnetization $m=\frac{1}{N}\sum_{i=0}^{i=N}s_{i}$. In each time step, one agent is selected at random, and outcomes of two types of dynamical rules performed independently (q-voter and q-Ising, described further) are checked (see Fig. 1). If both dynamics indicate a change of opinion, the selected node flips its state (a so-called AND rule [19]). Alternatively, one can describe the system as a duplex (two-layer multiplex network) with different dynamics on each level and the condition that the state of node has to be the same in both layers, which in the case of opinion dynamics, can be understood as “no hypocrisy rule”, i.e., the agent presents the same opinion in each community it belongs to.
@@
Figure 2: Top row: solutions of the rate equation (3) for (a) $q=1$, (b) $q=2$ and (c) $q=3$. The red solid line in the $m=0$ plane is $p^{*}_{q}(T)$ given by (5), while dashed lines come from the Taylor expansion thereof given by (6). Bottom row: comparison of numerical simulations and analytical predictions of $m(p)$ and $m(T)$ for (d,e) $q=1$, (f,g) $q=2$ and (h,i) $q=3$. Solid lines are solutions of $F_{1}=0$, $F_{2}=0$ and $F_{3}=0$ and symbols are Monte Carlo simulations for (d) $T=10$ – circles, $T=2$ – triangles, $T=1$ – diamonds, and $T=0.3$ – squares, (e) $p=0.99$ – circles, $p=0.5$ – upward triangles, $p=0.1$ –diamonds, $p=0.01$ – squares, and $p=0.001$ – downward triangles, (f) $T=100$ – circles, $T=2$ – triangles, $T=1$ – diamonds, $T=0.2$ – squares, (g) $p=0.71$ – circles, $p=0.68$ – upward triangles, $p=0.5$ –diamonds, $p=1/3$ – squares, and $p=0.3$ – downward triangles; (h) $T=1000$ – circles, $T=6$ – triangles, $T=3$ – diamonds, $T=1.73$ – squares; (i) $p=1$ – circles, $p=1/2$ – upward triangles, $p=1/3$ –diamonds, $p=0.2$ – squares. All simulations were performed for $N=2\cdot 10^{5}$ nodes and $M=2\cdot 10^{5}$ large MC steps.
Figure 3: Top row: solutions of the rate equation (3) for (a) $q=4$, (b) $q=5$ and (c) $q=6$. The red solid line in the $m=0$ plane is $p^{*}_{q}(T)$ given by (5), while dashed lines come from the Taylor expansion thereof given by (6). White solid lines denote the value of magnetization at the critical line $m(p^{*}_{q}(T))$. Bottom row: comparison of numerical simulations and analytical predictions of $m(p)$ and $m(T)$ for (d,e) $q=4$, (f,g) $q=5$, and (h,i) $q=6$. Solid lines are solutions of $F_{4}=0$, $F_{5}=0$ and $F_{6}=0$ and symbols are Monte Carlo simulations for (d) $T=100$ – circles, $T=10$ – up triangles, $T=5.5$ – diamonds, $T=3.5$ – squares, and $T=2.2$ – down triangles, (e) $p=1$ – circles, $p=0.8$ – up triangles, $p=0.4$ – diamonds, $p=3/11$ – squares, and $p=0.2$ – down triangles, (f) $T=10^{3}$ – circles, $T=7.5$ – triangles, $T=4$ – diamonds, $T=3.3$ – squares, (g) $p=1$ – circles, $p=0.4$ – up triangles, $p=0.25$ – diamonds, $p=0.2$ – squares, $p=0.15$ – down triangles; (h) $T=10^{3}$ – circles, $T=10$ – triangles, $T=5$ – diamonds, $T=3.9$ – squares; (i) $p=1$ – circles, $p=0.2$ – up triangles, $p=0.155$ – diamonds, $p=0.15$ – squares, $p=0.13$ – down triangles. All simulations were performed for $N=2\cdot 10^{5}$ nodes and $M=2\cdot 10^{5}$ large MC steps. Full symbols denote ordered initial conditions ($\langle m\rangle=1$), while empty symbols – random initial conditions ($\langle m\rangle=0$)
@@ III.1 The $q=1$ case
[…]
@@ III.2 The $q=2$ case
[…]
The last two solutions either bring no relevant values ($|m_{3,4}|>1$ or they are equal to ±$\pm 1$, but in such a case, they are unstable (see Appendix A for details). As for $q=1$, also here, the system exhibits a continuous phase transition for both control parameters (see analytical solutions of Eq. (11) in Fig. 2b and Monte Carlo simulations in Fig. 2f and in Fig. 2g). The relation between the critical value of $p$ and the temperature can once again be obtained using Eq. (5) or more directly by solving $A_{2}=0$ resulting in
@@ III.3 The $q=3$ case
and can be approximated by $p^{*}_{3}(T)=1/3+1/T$ for $T\gg 1$, closely covering the exact value (see solid and dashed lines in Fig. 2c). A notable difference between $q=3$ and $=1,2$ is connected to the fact that in addition to recovering the monolayer q-voter behavior for $T\rightarrow\infty$ (see 2h) when $p=1$, one should arrive at the original results of the monolayer q-Ising model. Below the critical temperature $T_{c}\approx 1.7214$ [11, 24], one observes no disordered state regardless of the value of $p$.
@@ IV.1 The $q=4$ case
The value of $T$ for which $B^{L}_{4}(T,p^{*}_{4}(T))$ changes its sign (provided that $C^{L}_{4}(T,p)>0$), indicates the tricrtical point that can be extracted as the only real, non-negative solution of $-55u^{4}-72u^{3}+358u^{2}+40u-15=0$ where $u=\exp(4/T_{\rm{tri}})$. Combined with $p^{*}_{4}(T_{\rm{tri}})$ leads to following coordinates of the tricritical point
@@ IV.2 The $q=5$ case
Judging from the properties gathered in Table 1, one expects that the coupled $q=5$ system should possess similar properties to the $q=4$ one. However, the system with $q=5$ can be characterized by a finite value of magnetization at the critical line even for large $T$ (cf. Fig. 3b,d,g, where we numerically calculated the magnetization for $m(T,p^{*}_{5}(T))$ along with numerical simulations). This result is also confirmed in the Landau approach (see Appendix C) by exploring $B(T,p)$ that approaches zero as $-T^{-1}$ for $T\gg 1$. Apparently, $m(T,p^{*}_{5}(T))$ is not a monotonic function and takes a maximum for $(T\approx 4.3207,p\approx 0.6309)$. As a result, we observe a narrow strip of bistability along the critical line that non-monotonically changes its width. These interesting differences between $q=4$ and $q=5$ can be attributed to the character of the phase transition for the base q-voter, i.e., a tricritical phase transition [10] where the magnetization at the critical point scales as $m\sim(p_{c}-p)^{\beta}$ with $\beta=1/4$ (unlike $q<5$ where $\beta=1/2$).
@@ IV.3 The $q=6$ case
Figure 5: Results for $q=20$. (a) Solution for the rate equation $F_{20}=0$ (b) Phase diagram for $q=20$. (c) Comparison of analytical results for $q=20$ and $p=0.39$ (solid line) and numerical simulations (symbols). Circles denote simulations starting from the disordered initial conditions $\langle m\rangle=0$, squares - the ones starting from the fully ordered state $\langle m\rangle=1$, and triangles - from $\langle m\rangle=0.4$. The dashed line marks the value of $T$ for which the effective potential was calculated, see panel (d). All simulations were performed for $N=10^{6}$ nodes and $M=2\cdot 10^{5}$ large MC steps. (d) Effective potential $V_{20}(m)$ for $p=0.39$ and $T=16.55$
@@ V Higher values of $q$
VI Conclusions
@@
The very construction of the coupled model gives the possibility to obtain the exact relation for the critical line $p^{*}_{q}(T)$, separating ordered and disordered states, virtually for any value of $q$, which takes a particularly simple form of $p^{*}_{q}(T)=\tanh T^{-1}$ if the lobby consists simple of one randomly chosen neighbor. Moreover, we also showed that in the high temperature limit, $p^{*}_{q}(T)$ can be rewritten as $T(p-a)=b$ with $a$ being the critical value of $p$ for the base q-voter model. Although only holding for the critical line, this relation suggests that temperature can be treated as a measure opposite to noise represented here by the independence probability $p$. This observation is most prominent for $q=1$, where in the high temperature limit we obtain $pT=1$.

### v3

@@ Abstract
We consider a coupled system mimicking opinion formation under the influence of a group of $q$ neighbors ($q$-lobby) that consists of an Ising part governed by temperature-like parameter $T$ and a voter dynamics parameterized by noise probability $p$ (independence of choice). Using rigorous analytical calculations backed by extensive Monte Carlo simulations, we examine the interplay between these two quantities. Based on the theory of phase transitions, we derive the relation between $T$ and $p$ at the critical line dividing the ordered and disordered phases, which takes a very simple and generic form T(pa)=bT(p-a)=b in the high temperature limit. For specific lobby sizes, we show where the temperature and noise are balanced, and we hint that for large $q$, the temperature-like dynamics prevails.
@@ I Introduction
Apart from exhibiting interesting statistical physics phenomena (e.g., first and second order phase transitions, tricriticality), both models in question are heavily influenced by social psychology research on the impact of conformity. In particular, we refer here to Solomon Asch experiments [19, 20] that have shown this impact to increase with the size of the unanimous influence group, but only up to a certain threshold; moreover, the conformity is reduced in a drastic manner if the group of influence is not unanimous. The probability $p$ in the q-voter model and the temperature-like parameter $T$ in its q-Ising counterpart can be used to tune the level of independence. However, individuals typically live simultaneously in different communities – be it offline or online – that can be treated as different levels in a multilayer network, where they are subject to opinion change suggestions. Such a setting has been investigated in the form of a duplex q-voter [21] and duplex q-Ising models [22] that have proven that the addition of a level dramatically changes the behavior of the system, e.g., while for the monolayer q-Ising we have second or first order phase transitions between the ordered and disorder phase depending on the lobby size, in the q-Ising duplex only second order transition takes place.
In this study, we couple these two different generic dynamics and we examine the conditions that are necessary for one of the models to prevail, thus finding the relation between the temperature and noise (independence). Additionally, owing to the parametrization of the model with the lobby size $q$, we show how the number of active neighbors affects the above relation and alters the system’s behavior.
@@ II Model description and its fundamental properties
The system consists of $N$ agents, each holding a binary opinion $s_{i}=\pm 1$, acting on top of a complete graph, described by total magnetization $m=\frac{1}{N}\sum_{i=0}^{i=N}s_{i}$ (average spin per node). In each time step, one agent is selected at random, and outcomes of two types of dynamical rules performed independently (q-voter and q-Ising, described further) are checked (see Fig. 1). If both dynamics indicate a change of opinion, the selected node flips its state – it is a so-called AND rule, previously used in threshold dynamics [23], majority-vote models [24], and q-voter [21, 25] and q-Ising dynamics [22]. Alternatively, one can describe the system as a duplex (two-layer multiplex network) with different dynamics on each level and the condition that the state of node has to be the same in both layers, which in the case of opinion dynamics, can be understood as “no hypocrisy rule”, i.e., the agent presents the same opinion in each community it belongs to. Different types of coupling may be considered, e.g., the OR rule, where it suffices to prompt the change on one level to flip the node. The AND rule has been chosen owing to the possibility of comparing with previous studies connected to duplex networks [21, 22, 26], but also to imitate the fact that acquiring the same opinion from more than one group reinforces it and allows it to last (in a similar way as it is done in the case of echo chamber effect for the spread of information [27, 28, 29]). In contrast, the OR rule should lead to fast suppression of the ordered phase (see the Discussion section for a detailed derivation of the rate equation for the OR rule and the description of differences between OR and AND rules).
@@
Figure 2: Top row: solutions of the rate equation (3) for (a) $q=1$, (b) $q=2$ and (c) $q=3$. The red solid line in the $m=0$ plane is $p^{*}_{q}(T)$ given by (5), while dashed lines come from the Taylor expansion thereof given by (6). Bottom row: comparison of numerical simulations and analytical predictions of $m(p)$ and $m(T)$ for (d,e) $q=1$, (f,g) $q=2$ and (h,i) $q=3$. Solid lines are solutions of $F_{1}=0$, $F_{2}=0$ and $F_{3}=0$ and symbols are Monte Carlo simulations for (d) $T=10$ – circles, $T=2$ – triangles, $T=1$ – diamonds, and $T=0.3$ – squares, (e) $p=0.99$ – circles, $p=0.5$ – upward triangles, $p=0.1$ –diamonds, $p=0.01$ – squares, and $p=0.001$ – downward triangles, (f) $T=100$ – circles, $T=2$ – triangles, $T=1$ – diamonds, $T=0.2$ – squares, (g) $p=0.71$ – circles, $p=0.68$ – upward triangles, $p=0.5$ –diamonds, $p=1/3$ – squares, and $p=0.3$ – downward triangles; (h) $T=1000$ – circles, $T=6$ – triangles, $T=3$ – diamonds, $T=1.73$ – squares; (i) $p=1$ – circles, $p=1/2$ – upward triangles, $p=1/3$ –diamonds, $p=0.2$ – squares. All simulations were performed for $N=2\cdot 10^{5}$ nodes and $M=2\cdot 10^{5}$ large MC steps [31].
Figure 3: Top row: solutions of the rate equation (3) for (a) $q=4$, (b) $q=5$ and (c) $q=6$. The red solid line in the $m=0$ plane is $p^{*}_{q}(T)$ given by (5), while dashed lines come from the Taylor expansion thereof given by (6). White solid lines denote the value of magnetization at the critical line $m(p^{*}_{q}(T))$. Bottom row: comparison of numerical simulations and analytical predictions of $m(p)$ and $m(T)$ for (d,e) $q=4$, (f,g) $q=5$, and (h,i) $q=6$. Solid lines are solutions of $F_{4}=0$, $F_{5}=0$ and $F_{6}=0$ and symbols are Monte Carlo simulations for (d) $T=100$ – circles, $T=10$ – up triangles, $T=5.5$ – diamonds, $T=3.5$ – squares, and $T=2.2$ – down triangles, (e) $p=1$ – circles, $p=0.8$ – up triangles, $p=0.4$ – diamonds, $p=3/11$ – squares, and $p=0.2$ – down triangles, (f) $T=10^{3}$ – circles, $T=7.5$ – triangles, $T=4$ – diamonds, $T=3.3$ – squares, (g) $p=1$ – circles, $p=0.4$ – up triangles, $p=0.25$ – diamonds, $p=0.2$ – squares, $p=0.15$ – down triangles; (h) $T=10^{3}$ – circles, $T=10$ – triangles, $T=5$ – diamonds, $T=3.9$ – squares; (i) $p=1$ – circles, $p=0.2$ – up triangles, $p=0.155$ – diamonds, $p=0.15$ – squares, $p=0.13$ – down triangles. All simulations were performed for $N=2\cdot 10^{5}$ nodes and $M=2\cdot 10^{5}$ large MC steps [31]. Full symbols denote ordered initial conditions ($\langle m\rangle=1$), while empty symbols – random initial conditions ($\langle m\rangle=0$)
@@ III.1 The $q=1$ case
$\left\{\begin{aligned} m_{0}~~&=~0\\ m_{1,2}&=\pm\sqrt{\frac{1-p\coth T^{-1}}{1-p}}\end{aligned}\right$. (8)
@@ III.2 The $q=2$ case
$\left\{\begin{aligned} m_{0}~&=~0\\ m_{1,2}&=\pm\sqrt{\frac{B_{2}+\sqrt{B_{2}^{2}-4A_{2}C_{2}}}{2C_{2}}}\\ m_{3,4}&=\pm\sqrt{\frac{B_{2}-\sqrt{B_{2}^{2}-4A_{2}C_{2}}}{2C_{2}}}\end{aligned}\right$. (12)
The last two solutions either bring no relevant values ($|m_{3,4}|>1$ or they are equal to $\pm 1$, but in such a case, they are unstable (see Appendix A for details). As for $q=1$, also here, the system exhibits a continuous phase transition for both control parameters (see analytical solutions of Eq. (11) in Fig. 2b and Monte Carlo simulations in Fig. 2f and in Fig. 2g). The relation between the critical value of $p$ and the temperature can once again be obtained using Eq. (5) or more directly by solving $A_{2}=0$ resulting in
@@ III.3 The $q=3$ case
and can be approximated by $p^{*}_{3}(T)=1/3+1/T$ for $T\gg 1$, closely covering the exact value (see solid and dashed lines in Fig. 2c). A notable difference between $q=3$ and =1,2=1,2 is connected to the fact that in addition to recovering the monolayer q-voter behavior for $T\rightarrow\infty$ (see 2h) when $p=1$, one should arrive at the original results of the monolayer q-Ising model. Below the critical temperature $T_{c}\approx 1.7214$ [11, 32], one observes no disordered state regardless of the value of $p$.
@@ IV.1 The $q=4$ case
The value of $T$ for which $B^{L}_{4}(T,p^{*}_{4}(T))$ changes its sign (provided that $C^{L}_{4}(T,p)>0$), indicates the tricrtical point that can be extracted as the only real, non-negative solution of $55u^{4}-72u^{3}+358u^{2}+40u-15=0$ where $u=\exp(4/T_{\rm{tri}})$. Combined with $p^{*}_{4}(T_{\rm{tri}})$ leads to following coordinates of the tricritical point
@@ IV.2 The $q=5$ case
Judging from the properties gathered in Table 1, one expects that the coupled $q=5$ system should possess similar properties to the $q=4$ one. However, the system with $q=5$ can be characterized by a finite value of magnetization at the critical line even for large $T$ (cf. Fig. 3b,d,g, where we numerically calculated the magnetization for $m(T,p^{*}_{5}(T))$ along with numerical simulations). This result is also confirmed in the Landau approach (see Appendix C) by exploring $B(T,p)$ that approaches zero as $T^{-1}$ for $T\gg 1$. Apparently, $m(T,p^{*}_{5}(T))$ is not a monotonic function and takes a maximum for $(T\approx 4.3207,p\approx 0.6309)$. As a result, we observe a narrow strip of bistability along the critical line that non-monotonically changes its width. These interesting differences between $q=4$ and $q=5$ can be attributed to the character of the phase transition for the base q-voter, i.e., a tricritical phase transition [10] where the magnetization at the critical point scales as $m\sim(p_{c}-p)^{\beta}$ with $\beta=1/4$ (unlike $q<5$ where $\beta=1/2$).
@@ IV.3 The $q=6$ case
Figure 5: Results for $q=20$. (a) Solution for the rate equation $F_{20}=0$ (b) Phase diagram for $q=20$. (c) Comparison of analytical results for $q=20$ and $p=0.39$ (solid line) and numerical simulations (symbols) [31]. Circles denote simulations starting from the disordered initial conditions $\langle m\rangle=0$, squares - the ones starting from the fully ordered state $\langle m\rangle=1$, and triangles - from $\langle m\rangle=0.4$. The dashed line marks the value of $T$ for which the effective potential was calculated, see panel (d). All simulations were performed for $N=10^{6}$ nodes and $M=2\cdot 10^{5}$ large MC steps. (d) Effective potential $V_{20}(m)$ for $p=0.39$ and $T=16.55$
@@ V Higher values of $q$
Figure 6: Higher values of $q$. (a) Log-log plot of the critical line for $p^{*}_{q}(T)$ given by Eq. (43) for $q=10$ (dark solid line), $q=20$ (light solid line), $q=30$ (dark dashed line), and $q=40$ (light dashed line), (b) Semi-log plot of critical values $p^{*}_{q}(T)$ versus $q$ given by Eq. (43) for $T=10$ (circles), $T=20$ (triangles), $T=30$ (diamonds), and $T=40$ (squares).
To examine the character of the lower spinodal in a more quantitative way, we derived a fully analytical formula for the critical line $p^{*}_{q}(T)$, shown separately as Eq. (43) in Appendix D due to its complicated form. This allows us to show in Fig. 6a that even for relatively small values of lobby size, the critical line runs almost vertically between $p=1$ and negligible values of $p$ at which it then stays regardless of the temperature. Although the limiting behavior for $q\to\infty$ is not traceable with Eq. (43) due to the mentioned linear relation between the critical temperature in the monolayer Ising model and the lobby size, we can still spot how the coupled system behaves at given $T$ with increasing $T$. Figure 6b suggests that the limiting case $p^{*}_{q}(T\to\infty)=2(q-1)/[(2(q-1)+2^{q}]$ which coincides with the monolayer q-voter lower spinodal (see Appendix D) serves as a satisfactory approximation.
VI Discussion
@@
The very construction of the coupled model gives the possibility to obtain the exact relation for the critical line $p^{*}_{q}(T)$, separating ordered and disordered states, virtually for any value of $q$, which takes a particularly simple form of $p^{*}_{q}(T)=\tanh T^{-1}$ if the lobby consists simple of one randomly chosen neighbor. Moreover, we also showed that in the high temperature limit, $p^{*}_{q}(T)$ can be rewritten as T(pa)=bT(p-a)=b with $a$ being the critical value of $p$ for the base q-voter model. Although only holding for the critical line, this relation suggests that temperature can be treated as a measure opposite to noise represented here by the independence probability $p$. This observation is most prominent for $q=1$, where in the high temperature limit we obtain $pT=1$.
@@
As mentioned in Sec. II, we have chosen the type of coupling (AND dynamics) motivated by its widespread use in the previous studies, providing an opportunity for a direct comparison. However, it is also possible to consider OR coupling, when the change is accepted if either of dynamics suggests a flip. From the user-centered point of view connected to multiplex networks, this rule brings to the front the fact that the opinion can be changed even if only one community (i.e., one layer) prompts for it. In such a situation the probability of changing a down spin to an up one is proportional to $\gamma^{+}_{I}(1-\gamma^{+}_{v})+\gamma^{+}_{v}(1-\gamma^{+}_{I})+\gamma^{+}_{I}\gamma^{+}_{v}$ (the spin flips if the q-Ising dynamics suggests a change and the q-voter one does not or if the q-voter dynamics suggests a change and the q-Ising one does not or if both do so). An opposite situation comes with $\gamma^{-}_{I}(1-\gamma^{-}_{v})+\gamma^{-}_{v}(1-\gamma^{-}_{I})+\gamma^{-}_{I}\gamma^{-}_{v}$ and as a result, the rate equation for the OR case reads
$F_{q}^{OR}=F^{I}_{q}+F^{v}_{q}-F^{AND}_{q}$ (18)
where $F^{I}_{q}=\frac{1}{2}[(1-m)\gamma^{+}_{I}-(1+m)\gamma^{-}_{I}]$ is the rate equation for the monolayer q-Ising model, $F^{v}_{q}=\frac{1}{2}[(1-m)\gamma^{+}_{v}-(1+m)\gamma^{-}_{v}]$ is the equivalent of the q-voter dynamics and $F^{AND}_{q}\equiv F_{q}$ given by Eq. (3) is the rate equation for the AND version examined in this study. The rate equation (18) proves that although OR dynamics is connected to the AND one, the base models play an important role that might suppress the coupling effect. Indeed, for the simplest case ($q=1$) one can easily obtain from (18) that $m_{0}=0$ and $m_{1,2}=\sqrt{(p+1)-(p-3)e^{-2/T}}/\sqrt{(1-p)\left(1-e^{-2/T}\right)}>1$, i.e., the system stays in the disordered state regardless of $p$ and $T$. This initial result is in line with the monolayer behavior of q-voter or q-voter dynamics for $q=1$ (see Table 1) while largely different than the system described by $F_{1}=0$ (see Fig. 2a) and also consistent with the lack of reinforcement introduced by the AND rule. However, these considerations are merely a glimpse into the problem of the OR dynamics that calls for a separate study that could examine it in detail.
In the same way, the presented model suffers from two obvious oversimplifications: the first one is connected to the limited number of opinions, while the second one is linked to the topology. Reaching outside the binary setting has solid psychological grounds, e.g., it gives the possibility to introduce an intermediate opinion ($s_{i}=0$) that could denote an unconvinced state. While the q-voter model has in fact been generalized for a higher number of states [39, 40], one would first need to find the equivalent of q-Ising version of the multi-state Ising model, i.e., the Potts model [41]. Additionally, as can be seen for the multi-state q-voter model [39, 40], even a simple shift to a three-state dynamics opens the field for different ways of accounting, e.g., for the anticonformity (e.g., taking any opinion other than the q-lobby [39] versus moving away from the q-lobby opinion [40]. Therefore, although a multi-state version of the problem considered in this study might be psychologically relevant, it cannot be achieved in a straightforward way.
As it concerns the underlying topology, the coupled model could be examined in a more real-world-like topologies, such as a heterogeneous networks with scale-free degree distribution $p(k)\sim k^{-\gamma}$, both numerically and analytically, using the pair-approximation (PA) approach introduced in the case of the q-voter in [42]. However, as shown by Jedrzejewski [42] and Gradowski and Krawiecki [25], the results for large average degree $\langle k\rangle$ overlap with the mean-field approach for the complete graph, for small $\langle k\rangle$ shifting the transition point close to $p=0$, On the other hand, Krawiecki and Gradowski have shown [30] that the introduction of heterogeneity to the q-Ising duplex with partial overlap [22] might change the character of the phase transition in specific cases. Therefore, making predictions on the behavior of our coupled system under large heterogeneity is difficult beforehand.